ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 131, 2002

Alejandre, M. J. and A. Ballester-Bolinches On a theorem of Berkovich

149

Aljadeff, E. and J. Sonn Fields for which the projective Schur subgroup is the whole Brauer group

375

Alon, N., M. Krivelevich and V. H. Vu On the concentration of eigenvalues of random symmetric matrices

259

Avila, A. and J. Bochi A formula with some applications to the theory of Lyapunov exponents

125

Ballester-Bolinches, A. See M. J. Alejandre

Bochi, J. See A. Avila

Bourgain, J. On the distribution of the Fourier spectrum of Boolean functions

269

Buzzi, J. and V. Maume-Deschamps Decay of correlations for piecewise invertible maps in higher dimensions

203

Cohen, A. M. and D. B. Wales Linearity of Artin groups of finite type

101

Cummings, J. and E. Schimmerling Indexed squares  61

Davidov, J. and O. Mus̆karov Harmonic almost-complex structures on twistor spaces

319

Estrada, B. and E. Martínez q-Trigonal Klein surfaces  361
Fan, A. How many intervals cover a point in Dvoretzky covering?  157

Fiebig, D., U.-R. Fiebig and M. Yuri Pressure and equilibrium states for countable state Markov shifts

221

Fiebig, U.-R. See D. Fiebig

Fujiwara, K. On the outer automorphism group of a hyperbolic group

277

Geschke, S., M. Kojman, W. Kubiś and R. Schipperus Convex decompositions in the plane and continuous pair colorings of the irrationals

285

Kalenda, O. F. K. A new Banach space with Valdivia dual unit ball 139

Kojman, M. See S. Geschke

Krivelevich, M. See N. Alon

Kubiś, W. See S. Geschke

Ma, X. Functoriality of real analytic torsion forms  1

Martínez, E. See B. Estrada Maume-Deschamps, V. See J. Buzzi Montgomery-Smith, S. Rearrangement invariant norms of

symmetric sequence norms of independent sequences of random variables

51

Mus̆karov, O. See J. Davidov Rodríguez, J. L., J. Scherer and J. Thévenaz Finite simple

groups and localization

185

Scheerer, H., D. Stanley and D. Tanré Fibrewise construction applied to Lusternik–Schnirelmann category

333

Scherer, J. See J. L. Rodríguez Schimmerling, E. See J. Cummings Schipperus, R. See S. Geschke Sonn, J. See E. Aljadeff Stanley, D. See H. Scheerer Tanré, D. See H. Scheerer Thévenaz, J. See J. L. Rodríguez Vu, V. H. See N. Alon Wales, D. B. See A. M. Cohen Yuri, M. See D. Fiebig